OpenStudy (anonymous):

In a triangle ABC, angle C is 3 times as large as angle A. Angle B measures 5 degrees less than angle A. Find the measure of each angle. Talk me through it, please.

OpenStudy (anonymous):

Lets call angle A 'x'. Because angle C is "3 times as large" as A, it will be '3x' angle B will be 'x-5' because it is "5 degrees less than A" The sum of the angles must be 180, so we have: x+(3x)+(x-5) = 180 Simplifying this gives: 5x-5=180 5x=185 x=37 so angle A is 37, B is 37-5 = 32, and C is 3(37) = 111

OpenStudy (anonymous):

Thank you so much!!!! It has been many many years since I did this kind of stuff.